OpenStudy (anonymous):

HELP PLEASE!!! :( Suppose the probability of a major earthquake on a given day is 1 out of 15,000. use the poisson distribution to approximate the probability that there will be at least one major earthquake in the next 2000 days.

OpenStudy (kropot72):

\[\lambda=\frac{1}{15000}\times 2000=\frac{2}{15}\] \[P(X=1)=\frac{e ^{-\lambda}\lambda ^{1}}{1!}=you\ can\ calculate\] Then you will need to calculate P(X = 2), P(X = 3) and so on until the value becomes insignificant. After that the appropriate values of probability must be added to find the probability that there will be at least one major earthquake in the next 2000 days.

OpenStudy (anonymous):

thank you very much!

OpenStudy (kropot72):

You're welcome :)